# Reinforcement Learning for Causal Discovery without Acyclicity Constraints

### v3

@@ Abstract
Recently, reinforcement learning (RL) has proved a promising alternative for conventional local heuristics in score-based approaches to learning directed acyclic causal graphs (DAGs) from observational data. However, the intricate acyclicity constraint still challenges the efficient exploration of the vast space of DAGs in existing methods. In this study, we introduce $\mathbf{\mathrm{\mathbf{ALIAS}}}$ (reinforced dAg Learning wIthout Acyclicity conStraints), a novel approach to causal discovery powered by the RL machinery. Our method features an efficient policy for generating DAGs in just a single step with an optimal quadratic complexity, fueled by a novel parametrization of DAGs that directly translates a continuous space to the space of all DAGs, bypassing the need for explicitly enforcing acyclicity constraints. This approach enables us to navigate the search space more effectively by utilizing policy gradient methods and established scoring functions. In addition, we provide compelling empirical evidence for the strong performance of $\mathbf{\mathrm{\mathbf{ALIAS}}}$ in comparison with state-of-the-arts in causal discovery over increasingly difficult experiment conditions on both synthetic and real datasets.
@@ 5.1 Motivation for Reinforcement Learning
Using the $\mathbf{\mathrm{\mathbf{Vec2DAG}}}$ representation, the score-based causal discovery problem may seem to simplify into a maximization problem: $\mathbf{z}^{*}=\underset{\mathbf{z}\in\mathbb{R}^{d\cdot\left(d+1\right)/2}}{\arg\max}\mathcal{S}_{\text{BIC}}\left(\mathcal{D},\mathbf{\mathrm{\mathbf{Vec2DAG}}}\left(\mathbf{z}\right)\right)$, which could, in principle, be addressed to certain extents using off-the-shelf black-box optimization techniques such as Bayesian optimization, which is one of the most popular blackbox optimization methods. However, solving this optimization problem is far from straightforward due to the high dimensionality of the search space, which grows quadratically with the number of nodes (e.g., for 30 nodes, the space is 465-dimension). Meanwhile, for example, Bayesian optimization is typically limited to only tens of dimensions (Malu et al., 2021). Moreover, the number of optimization steps also poses a significant challenge, particularly for Bayesian optimization, which typically scales cubically with the number of steps.
Since RL is the only black-box optimization approach that has been studied in the score-based causal discovery literature (up to when our manuscript is written, to the best of our knowledge), we align with the established line of works (Zhu et al., 2020; Wang et al., 2021; Yang et al., 2023a; b) to specifically focus on leveraging RL to solve this optimization problem. The idea is that an RL agent with a stochastic policy can autonomously decide where to explore based on the uncertainty of the learned policy, which is continuously updated through incoming reward signals (Zhu et al., 2020). As shown below, RL provides our method with built-in exploration-exploitation capabilities and linear scalability with respect to both dimensionality and sample size, making it a practical choice for this problem. In addition, our RL point of view also enables the adaptability of various RL methods, such as vanilla policy gradient (Sutton et al., 1999), A2C (Mnih et al., 2016), and PPO (Schulman et al., 2017), to effectively optimize our objective, effectively establishing a clear association between our proposed approach and RL.
@@ 5.2 Policy Gradient for DAG Search
Utilizing RL, we seek for a policy $\pi$ that outputs a continuous action $\mathbf{z}\in\mathbb{S}_{d}=\mathbb{R}^{d\cdot\left(d+1\right)/2}$, which is the parameter space of DAGs of $d$ nodes. In this work, we consider stochastic policies for better exploration, i.e., we parametrize our policy by an isotropic Gaussian distribution with learnable means and variances: $\pi_{\bm{\theta}}\left(\mathbf{z}\right)=\mathcal{N}\left(\mathbf{z};\bm{\mu}_{\bm{\theta}},\text{diag}\left(\mathbf{\bm{\sigma}}_{\bm{\theta}}^{2}\right)\right)$. Since our policy generates a DAG representation in just one step, every trajectory starts with the same initial state and terminates after only one transition, so the agent does not need to be aware of the state in our method. We note that, similar to RL-BIC (Zhu et al., 2020), the one-step nature of the environment does not preclude the application of RL in our approach. This is because a one-step environment is simply a special case of a Markov decision process (MDP), which remains compatible with most RL algorithms.

### v4

@@ Abstract
Recently, reinforcement learning (RL) has proved a promising alternative for conventional local heuristics in score-based approaches to learning directed acyclic causal graphs (DAGs) from observational data. However, the intricate acyclicity constraint still challenges the efficient exploration of the vast space of DAGs in existing methods. In this study, we introduce $\mathbf{\mathrm{\mathbf{ALIAS}}}$ (reinforced dAg Learning wIthout Acyclicity conStraints), a novel approach to causal discovery powered by the RL machinery. Our method features an efficient policy for generating DAGs in just a single step with an optimal quadratic complexity, fueled by a novel parametrization of DAGs that directly translates a continuous space to the space of all DAGs, bypassing the need for explicitly enforcing acyclicity constraints. This approach enables us to navigate the search space more effectively by utilizing policy gradient methods and established scoring functions. In addition, we provide compelling empirical evidence for the strong performance of $\mathbf{\mathrm{\mathbf{ALIAS}}}$ in comparison with state-of-the-arts in causal discovery over increasingly difficult experiment conditions on both synthetic and real datasets. Our implementation is provided at https://github.com/baosws/ALIAS.
@@ 5.1 Motivation for Reinforcement Learning
Using the $\mathbf{\mathrm{\mathbf{Vec2DAG}}}$ representation, the score-based causal discovery problem may seem to simplify into a maximization problem: $\mathbf{z}^{*}=\underset{\mathbf{z}\in\mathbb{R}^{d\cdot\left(d+1\right)/2}}{\arg\max}\mathcal{S}_{\text{BIC}}\left(\mathcal{D},\mathbf{\mathrm{\mathbf{Vec2DAG}}}\left(\mathbf{z}\right)\right)$, which could, in principle, be addressed to certain extents using off-the-shelf black-box optimization techniques such as Bayesian optimization, which is one of the most popular blackbox optimization methods. However, solving this optimization problem is far from straightforward due to the high dimensionality of the search space, which grows quadratically with the number of nodes (e.g., for 30 nodes, the space is 465-dimension).
Since RL is the only black-box optimization approach that has been studied in the score-based causal discovery literature (at the time writing this manuscript, to the best of our knowledge), we align with the established line of works (Zhu et al., 2020; Wang et al., 2021; Yang et al., 2023a; b) to specifically focus on leveraging RL to solve this optimization problem. The idea is that an RL agent with a stochastic policy can autonomously decide where to explore based on the uncertainty of the learned policy, which is continuously updated through incoming reward signals (Zhu et al., 2020). That said, we note that, our method, which only involves one-step trajectories as shown below, is better perceived as a policy gradient approach rather than a general RL method that requires a multi-step Markov decision process (MDP).
As shown below, policy gradient provides our method with built-in exploration-exploitation capabilities and linear scalability with respect to both dimensionality and sample size, making it a practical choice for this problem. In addition, our policy gradient point of view also enables the adaptability of various established methods, such as vanilla policy gradient (Sutton et al., 1999), A2C (Mnih et al., 2016), and PPO (Schulman et al., 2017), to efficiently optimize our objective, effectively establishing a clear association between our proposed approach and policy gradient in RL.
@@ 5.2 Policy Gradient for DAG Search
Utilizing RL, we seek for a policy $\pi$ that outputs a continuous action $\mathbf{z}\in\mathbb{S}_{d}=\mathbb{R}^{d\cdot\left(d+1\right)/2}$, which is the parameter space of DAGs of $d$ nodes. In this work, we consider stochastic policies for better exploration, i.e., we parametrize our policy by an isotropic Gaussian distribution with learnable means and variances: $\pi_{\bm{\theta}}\left(\mathbf{z}\right)=\mathcal{N}\left(\mathbf{z};\bm{\mu}_{\bm{\theta}},\text{diag}\left(\mathbf{\bm{\sigma}}_{\bm{\theta}}^{2}\right)\right)$. Since our policy generates a DAG representation in just one step, every trajectory starts with the same initial state and terminates after only one transition, so the agent does not need to be aware of the state in our method. We note that, similar to RL-BIC (Zhu et al., 2020), the one-step nature of the environment does not preclude the application of RL in our approach. This is because a one-step environment is simply a special case of a MDP, which remains compatible with most RL algorithms.
@@ D.7 Comparison with Constraint-based methods
D.8 On Varsortability
Recently, it was suggested by Reisach et al. (2021) that marginal variances share high dependencies with the topological ordering of the true DAG for simple additive noise models. As such, sorting the variables according to the variances may be sufficient to recover the DAG, voiding the need for more sophisticated methods. More particularly, the metric $\texttt{Varsortability}\in\left[0,1\right]$ is proposed to assess the correlation between the ordering of marginal data variances and the causal ordering of the true DAG. Essentially, Varsortability calculates the portion of directed paths in the ground-truth DAG that have increasing marginal variances, and thus, if it is close to 1 then there is a high chance that simply sorting the variances would reveal the true DAG, and vice versa. It was shown that simple additive noise data commonly employed in existing studies, starting with Zheng et al. (2018), exhibit high values of Varsortability.
However, we show here that this is not entirely the case for our study. Particularly, while the majority of existing studies employ sparse graphs for evaluation, where the densest graphs have an expected in-degree of about 4 at most (Zheng et al., 2018; Yu et al., 2021; Bello et al., 2022; Massidda et al., 2024), in this study, we also consider much denser graphs with an expected in-degree of up to 8. This distinction is of high importance because we find that Varsortability decreases rapidly with the graph density and reaches near-zero at ER-8 graphs (Table 24), indicating that our data settings is much more nontrivial, and cannot be resolved simply by sorting the variances. Furthermore, since Varsortability is close to 0 in our data, one may expect that simply sorting the nodes by decreasing variances may work. In Table 25 we show that this is not the case, where the sortnregress algorithm proposed in conjunction to Varsortability (Reisach et al., 2021)—which first sorts the nodes by marginal variances and then perform linear feature selection to convert the ordering to a DAG—cannot recover the true DAG regardless of the sorting direction. Nevertheless, under this intricate scenarios, our method still robustly recovers the true DAG with a near-zero SHD.
Table 24: Varsortability as function of graph density. We compute Varsortability (Reisach et al., 2021) for linear-Gaussian data on 30-node ER graphs (Section. 6.2) with varying expected in-degrees. The numbers are mean ± standard errors over 100 simulations.
Expected in-degree Varsortability
1 $0.97\pm 0.01$
2 $0.87\pm 0.03$
4 $0.35\pm 0.06$
6 $0.05\pm 0.01$
8 $0.01\pm 0.00$
Table 25: Comparison with sortnregress on dense data. We consider linear-Gaussian datasets on dense 30-node ER-8 graphs. The numbers are mean ± standard errors over 5 simulations.
Method SHD ($\downarrow$) FDR ($\downarrow$) TPR ($\uparrow$)
sortnregress (decreasing variances) $360.2\pm\phantom{0}2.2$ 0.97±0.01 0.04±0.01
sortnregress (increasing variances) $\phantom{0}88.2\pm 17.7$ 0.27±0.05 0.94±0.01
$\mathbf{\mathrm{\mathbf{ALIAS}}}$ (ours) $\mathbf{\phantom{0}\phantom{0}0.2\pm\phantom{0}0.2}$ $\mathbf{0.00\pm 0.00}$ $\mathbf{1.00\pm 0.00}$
D.9 Effect of Data Standardization
Given the issues raised by Reisach et al. (2021) as discussed in the previous section, it is interesting to study the performance of causal discovery methods under data standardization, where the variance ordering information is completely removed.
Linear-Gaussian data. For linear-Gaussian data, the causal graph is identifiable under the equal-noise variance assumption (Peters et al., 2014). Hence, standardizing our linear-Gaussian data node-wise would render the noise variances unequal and reduce the causal model to a general linear-Gaussian model, where the causal graph is known to be unrecoverable (Spirtes et al., 2000). As such, it is expected that causal discovery performance on standardized linear data is worse than its non-standardized counterparts. However, in Table. 26 we show that our $\mathbf{\mathrm{\mathbf{ALIAS}}}$ method can still significantly outperform the baselines under this setting with the lowest SHD, which is only half of that of the second-best method, highlighting $\mathbf{\mathrm{\mathbf{ALIAS}}}$’s robustness against data standardization.
Nonlinear data. On the contrary with linear causal models, nonlinear additive noise models are generally identifiable regardless of unequal noise variances (Hoyer et al., 2008), meaning standardizing data should have minimal influence on the causal discovery performance. Indeed, in Table. 26 we show that our $\mathbf{\mathrm{\mathbf{ALIAS}}}$ method still performs equally well when the data is standardized compared to when the data is not standardized (Table. 4) with a very low SHD of only around 1, which is much lower relative to other baselines.
Table 26: Causal Discovery performance on standardized data. We consider linear-Gaussian data and nonlinear data with Gaussian processes. Observed data is standardized dimension-wise to remove variance information. For $\mathbf{\mathrm{\mathbf{ALIAS}}}$, we use the non-equal variances BIC version to reflect the data condition. The numbers are mean ± standard errors over 5 independent runs. Bold: best performance, underline: second-best performance.
Linear data (10-node ER-2 graphs) Nonlinear data (10-node ER-4 graphs)
NOTEARS (Zheng et al., 2020) $\underline{10.2\pm 2.0}$ $\underline{0.39\pm 0.04}$ $\mathbf{0.84\pm 0.02}$ $7.8\pm 2.3$ $0.14\pm 0.07$ $0.91\pm 0.01$
DAGMA (Bello et al., 2022) $17.4\pm 2.8$ $0.70\pm 0.08$ $0.39\pm 0.07$ $29.8\pm 5.5$ $0.4\pm 0.16$ $0.28\pm 0.15$
COSMO (Massidda et al., 2024) $16.4\pm 1.9$ $0.74\pm 0.06$ $0.29\pm 0.06$ $30.8\pm 4.9$ $0.5\pm 0.16$ $0.27\pm 0.12$
$\mathbf{\mathrm{\mathbf{ALIAS}}}$ (ours) $\mathbf{\phantom{0}6.0\pm 1.8}$ $\mathbf{0.30\pm 0.07}$ $\underline{0.79\pm 0.05}$ $\mathbf{1.4\pm 0.9}$ $\mathbf{0.0\pm 0.01}$ $\mathbf{0.97\pm 0.03}$